OpenStudy (anonymous):

Write the equation of a line in slope intercept form that passes through (2, 4) and (5, 4). the 2 slopes make a horozontal line. and I know that it would be 4-4=slope(5-2) i think. but would the slope be 0?

OpenStudy (anonymous):

@Hero Help?

hero (hero):

Yes, you are correct. The slope is zero. Whenever the y coordinates of two points are the same, the slope is zero.

OpenStudy (anonymous):

Here are some examples... sometimes helps you if you can learn to visualize what you are trying to solve with numbers: 

hero (hero):

@JakeV8, don't confuse her with all those lines bro

OpenStudy (anonymous):

ok, point taken :) @shannon.janette1, as I told you earlier, you have good help with @Hero... disregard my sketch if it confuses rather than helps.

OpenStudy (anonymous):

lol! Thanks to both of you! The sketch will help me with later problems! But it would be 4-4=0(5-2) correct?

OpenStudy (anonymous):

the way you wrote that is a little confusing, but I can tell you're saying that slope is 0, right? :) That is correct.

OpenStudy (anonymous):

How should I write it so that it isn't so confusing?

OpenStudy (anonymous):

slope = rise/run = (y2 - y1) / (x2 - x1) = (4-4)/(5-2) = 0 This is equivalent to what you wrote, though :)

OpenStudy (anonymous):

If I do write it like that woiuld my teacher still understand it?

OpenStudy (anonymous):

I would think so... you start with (y2 - y1) = m(x2 - x1) and then plug in the points (x1, y1) and (x2, y2) and solve for the slope m. If you show that work (which you did, mostly, up earlier in this problem), you should be fine.

hero (hero):

Make sure you write your fractions in a top-down manner @shannon.janette1

OpenStudy (anonymous):

what do you mean by "top down manner"? You mean instead of like a/b, show them in a normal fraction listing like:\[\frac{ a }{ b }\]